Are these parallel, perpendicular or neither
y=4x-1/3
2y=-1/2x+3

Answer by mananth(16946)

y=4x-1/3
2y=-1/2x+3
y= (-1/4)x+3
compare the slopes
4*-1/4 = -1
m1*m2 =-1
so they are perpendicular